OpenStudy (anonymous):

Standard form of a quadratic function what is the vertex form of the following function? y= 2/3x^2-8/3x+5/3 ~I know I can do this so I don't need it to be explained. please show all work and answer

OpenStudy (ranga):

y= 2/3x^2-8/3x+5/3 3y = 2x^2 - 8x + 5 3/2y = x^2 - 4x + 5/2 = (x-2)^2 - 4 + 5/2 = (x-2)^2 - 3/2 y = 2/3(x-2)^2 - 1

OpenStudy (anonymous):

thank you

OpenStudy (ranga):

you are welcome.